A roofer and an assistant can repair a roof together in 4 hours. Working alone, the assistant can complete the repair in 13 hours. If both the roofer and the assistant work together for 3 hours and then the assistant is left alone to finish the job, how much longer should the assistant need to finish the repairs? (Round your answer to one decimal place.)
Answer by lwsshak3(11628)

A roofer and an assistant can repair a roof together in 4 hours. Working alone, the assistant can complete the repair in 13 hours. If both the roofer and the assistant work together for 3 hours and then the assistant is left alone to finish the job, how much longer should the assistant need to finish the repairs? (Round your answer to one decimal place.)
***
1/13=work rate of assistant
1/4=work rate for roofer and assistant working together
..
3(1/4)+x/13=1
3/4+x/13=1
x/13=1/4
x=13/4=3.4
how much longer should the assistant need to finish the repairs? 3.4 hrs